Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2018-10-11

Observers:  T. Reynolds, S.Moran
Support Team: H. Kuncarayakti
TAT: Régis Cartier, Michel Dennefeld

Calibrations: 
Flats and arcs for gr11,13,16,18,20 + biases

Weather: 
Start - Thick clouds (closed)
Middle - Thick clouds, sleet (closed)
End - Cirrus, clearing

UT       Target       Type      Setup          Exp         airmass  DIMM  note
-----------------------------------------------------------------------------------------
                                             
EFOSC          
08:20 Opened
08:30    AT2018hee        Class                gr13+1"                600                        1.4          -        ia around max   
08:45    AT2018gle         Class                gr13+1"                300                        1.05        1.4         Ic?
09:00     AT2018hfr        Class                gr13+1"                300                            2         1.35        Ia-91T, 1 week pre-max   
09:10    AT2018hdq        Class                gr13+1"               600                          1.5        1.4        Not sure transient was in slit                     

Comments